OpenStudy (anonymous):

Help and explain? Please? A pendulum L inches in length takes t seconds to make one full cycle according to the equation t = 2π√L / 384 To the nearest tenth of an inch, what is the length of a pendulum that completes one full cycle every 1.5 seconds?

OpenStudy (anonymous):

The easiest way to do this type of problem is to first rearrange the equation with the variable you want on one side, and everything else on the other, so in this case you're going to want to get the variable L by its self!

OpenStudy (anonymous):

How do I get L by itself? It's in the square root sign.

OpenStudy (anonymous):

Well first can you get everything else to the other side so you just have sqrt of L?

OpenStudy (anonymous):

Okay, so first do I get rid of the 384? I think you times it by T.

OpenStudy (anonymous):

Right! So you've gotten that to the other side, now how would you move the 2 pi

OpenStudy (anonymous):

Minus 2 pi?

OpenStudy (anonymous):

Not quite! Because the sqrt t and 2 pi are multiplied, you would have divide by two pi on both sides of the equation! \[\sqrt{L} = \frac{384 t}{2 \pi}\] Now, how to get rid of the square root? What's the opposite of it?

OpenStudy (anonymous):

Oh, I understand now! So would you square both sides?

OpenStudy (anonymous):

Exactly! \[L = (\frac{384 t}{2 \pi })^{2}\] Now just plug in the value of t, and plug and chug!

OpenStudy (anonymous):

I plug t in and I'm not getting the right answer. Am I doing something wrong?

OpenStudy (anonymous):

T would be 1.5.

OpenStudy (anonymous):

\[L = (\frac{384 (1.5)}{2 \pi})^{2} = 8404.0\]

OpenStudy (anonymous):

Same thing I got!

OpenStudy (anonymous):

Sadly, none of the answer choices. :(

OpenStudy (anonymous):

(21.9) (9.6) (14.6) or (29.2)? I wonder what I did wrong.

OpenStudy (anonymous):

OH it said to the "nearest tenth of an inch" does that have something to do with it?

OpenStudy (anonymous):

Was the 384 also under a square root?

OpenStudy (anonymous):

Yes, 384 and L.

OpenStudy (anonymous):

Oh! I'm sorry, well then that changes it a bit! I'm sorry! I didn't think they would have given you it as a constant under it...so, we would have done all of the same things, when we moved the 384 it was under a square root so the equation was really... \[L = (\frac{\sqrt{384} t}{2 \pi})^{2}\] but only the 384 is under the sqare root! this should you give you the answer 21.9 approximately!

OpenStudy (anonymous):

Ohhhh, sorry for not telling you! It's fine, and I understand it now. Thanks you helped alot! :)